. A bicycle has wheels with a diameter of 42cm.

The bicycle is ridden in a straight line at a constant speed. The wheel makes 250

revolutions per minute.

What is the speed of the bicycle in kilometres per hour?


1
Expert's answer
2021-05-11T07:24:26-0400
"v=\omega r\\v=2\pi(250)(0.21)\frac{60}{1000}=19.8\frac{km}{h}"